OpenStudy (zned6559):

Calculate lim t-> negative infinity ((2t^2 - t -2)^1/2) / (2t+1)

OpenStudy (amistre64):

x = sqrt(x^2) ... might be useful

OpenStudy (zned6559):

I know the answer is -1/sqrt 2 but I don't know how to get there

OpenStudy (amistre64):

^(1/2) is the same as sqrt() id put them both under the sqrt and compare them

OpenStudy (amistre64):

\[\frac{\sqrt{2t^2-t-1}}{\sqrt{(2t+1)^2}}\]

OpenStudy (zned6559):

the bottom is suppose to be 2t+1 no sqrt

OpenStudy (amistre64):

what does sqrt(x^2) reduce to?

OpenStudy (zned6559):

x

OpenStudy (amistre64):

then sqrt((2t+1)^2) = 2t+1

OpenStudy (zned6559):

\[\sqrt{2t ^{2}-t-2}\div2t+1\]

OpenStudy (amistre64):

you are not seeing the obvious ... \[\sqrt{2t ^{2}-t-2}\div2t+1\] \[\sqrt{2t ^{2}-t-2}\div \sqrt{(2t+1)^2}\]

OpenStudy (amistre64):

what is the expansion of (2t+1)^2 ?

OpenStudy (zned6559):

\[4t ^{2}+4t+1\]

OpenStudy (amistre64):

then we have \[\sqrt{\frac{2t^2-t-2}{4t^2+4t+1}}\]

OpenStudy (amistre64):

what do we know about polynomials and their limits?

OpenStudy (zned6559):

not sure

OpenStudy (amistre64):

then you need to cover some of your basic rules .... not sure what theyve covered tho

OpenStudy (amistre64):

polys of the same degree, limit to their leading coeffs

OpenStudy (zned6559):

I got it now!

OpenStudy (amistre64):

the only thing to watch for is the sign, by squaring the bottom we altered the sign of it inthe process.

OpenStudy (amistre64):

2t+1 as t to -inf is a negative value .... so we are originally going to have a negative limit in the end